OpenStudy (anonymous):

What is the derivative of y= (xlnx)/(1+lnx)?

OpenStudy (tkhunny):

Well, you have some Product Rule and Quotient Rule in your future. Let's see your best efforts.

OpenStudy (anonymous):

\[ y =\frac{x \cdot ln(x)}{1 + ln(x)} \] \[ f(x) = x \cdot ln(x) \\ f'(x) = \bigg[ \; x \cdot ln(x) \; \bigg]' = (x)' \cdot ln(x) + x \cdot \big[ \; ln(x) \; \big]' =\\ = 1 \cdot ln(x) + x \cdot \frac{1}{x} = ln(x) + 1 \] \[ g(x) = 1 + ln(x) \\ g'(x) = \bigg[ \; 1 + ln(x) \; \bigg]' = (1)' + \big[\; ln(x) \; \big]' =0 +\frac{1}{x} = \frac{1}{x} = x^{-1} \\ \] \[ y =\frac{x \cdot ln(x)}{1 + ln(x)} = \frac{f(x)}{g(x)}\\ y' = \bigg[\frac{f(x)}{g(x)} \bigg]' = \frac{f'(x)g(x) - f(x)g'(x)}{\big[g(x)\big]^2} = \\ = \frac{ (ln(x) + 1)(1 + ln(x)) - x\cdot ln(x) \cdot x^{-1} } { \big[ \; ln(x) + 1 \; \big]^2 }=\\ = 1 - \frac{ ln(x) } { \big[ \; ln(x) + 1 \; \big]^2 } \] If I got it right..

OpenStudy (anonymous):

Seems right: http://www.wolframalpha.com/input/?i=%28x*ln%28x%29%29%2F%28ln%28x%29+%2B+1%29 Since \[ 1 - \frac{ln(x)}{\big[\;ln(x) + 1\;\big]^2 } = \frac{\big[\;ln(x) + 1\;\big]^2 -ln(x)}{\big[\;ln(x) + 1\;\big]^2 } = \\ = \frac{ln^2(x) + 2ln(x) + 1 -ln(x)}{\big[\;ln(x) + 1\;\big]^2 } = \\ = \frac{ln^2(x) + ln(x) + 1}{\big[\;ln(x) + 1\;\big]^2 } \] Which is what's in there